OpenStudy (mayaal):

find the arctan of 1 without a calculator

OpenStudy (aum):



OpenStudy (aum):

tan(t) = opposite / adjacent = 1 / 1 = 1 ABC is a right triangle with angle B = 90 degrees. AB = BC (isosceles triangle) Therefore, angle A = angle C. angles A + B + C = 180 A + 90 + A = 180 2A = 90 A = 45 degrees So arctan(1) = 45 degrees.

OpenStudy (mayaal):

i just wanted to make sure if its pi/4.

OpenStudy (aum):

Yes. pi/4 radians or 45 degrees.

OpenStudy (mayaal):

ok,thnku very much 4 helping me once again:)

OpenStudy (aum):

you are welcome.